# Socialized Learning: A Survey of the Paradigm Shift for Edge Intelligence in Networked Systems

### v1

@@ V-F1 Unidirectional Competition - Compulsory
Compulsory interaction is characterized by the dominance of one group and the subsequent compliance of another. Often, this stems from structural hierarchies, such as in workplaces where managers direct employees, or in social constructs where certain groups hold sway over others. In ML, this interaction can be observed in hierarchical structures where higher-level algorithms or protocols dictate the functioning or constraints of those at lower levels. By understanding this unidirectional competition, researchers can harness the efficiency and directionality it offers, while are mindful of the inherent constraints and potential lack of flexibility it introduces.
@@
HRL mirrors unidirectional compulsory interactions where a dominant group or layer prescribes rules or guidelines for another group, mandating compliance. This structure is visually represented in Fig. 21. The essence of compulsory directionality is exhibited as the higher echelons set overarching goals, while the subordinate layers are obligated to adhere. Dayan et al. [153] offer a quintessential example with a system partitioned into managers and sub-managers. Governed by principles like reward hiding and information hiding, this structure encapsulates the unidirectional competition, where each layer, while autonomous to some extent, is invariably tethered to the superiors.
Figure 21: Illustration of HRL in Socialized Training.
@@
where $H$ is the main function, $s$ is the state, $\pi^{G}$ and $\pi^{g}$ are the high-level and low-level policies, $\text{Transit}(\cdot)$ is the transition function, and $R$ is the reward function. Such hierarchical frameworks not only streamline learning but also afford the assimilation of prior knowledge. As in social structures, HRL embodies the edicts of one group directing the behavior of another. The directives from higher layers, while potentially restrictive, shape the trajectory of the entire system, ensuring alignment with overarching objectives [154].
@@ V-F2 Bidirectional Competition - Struggle
The concept of struggle, in social dynamics when entities vie for shared resources, power, or influence, is reflected in adversarial learning techniques. In these techniques, two components challenge each other, and it’s through this continuous push and pull that the model evolves, yielding improved outcomes.
V-F2a Adversarial Learning
In adversarial learning, particularly notable in generative adversarial networks (GANs) [155], there is a struggle between two components: the generator and the discriminator, as illustrated in Fig. 22. The generator crafts authentic samples to deceive the discriminator. Conversely, the ever-vigilant discriminator endeavors to discern real data from the counterfeits created by the generator. However, challenges like instability in the training process and homogeneous sample generation persist. Original GANs often produced uncontrollable image variations from random noise.
[…]
$\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))],$
@@
Solutions emerged in the form of Conditional GANs [156], introducing conditional data, and DCGANs [157], swapping fully connected layers for convolutional ones. The quest for stability birthed the Wasserstein GAN [158] with its novel loss function. Drawing parallels to social struggle, the adversarial learning framework represents inter-layer conflicts. This dynamic framework extends beyond image generation, spanning areas like RL [159] and NLP [160].
Communication-Aware Adversarial Learning: Adversarial learning techniques, such as GANs, have shown remarkable success in various applications. However, in a distributed setting, the communication between the generator and discriminator networks can be costly and may impact the overall performance. Communication-aware adversarial learning techniques have been proposed to address this issue by reducing the communication overhead while maintaining the quality of the generated samples. These methods may involve strategies such as gradient compression, selective parameter sharing, and decentralized architectures for adversarial learning. The design of communication-efficient adversarial learning systems often requires considering the trade-off between communication cost and the stability and convergence of the training process [161].
@@
Decentralized Learning [86, 23, 87, 88, 89] Peer-to-peer collaboration among devices with limited resources. Efficient resource utilization, improved privacy, and robustness against single points of failure. Increased communication overhead, potential for slower convergence.
@@
Meta Learning [132, 133] Applications requiring fast adaptation to new tasks or environments. Enables rapid learning and adaptation to new tasks, improves model generalization. Increased computational complexity, challenges in defining appropriate meta-learning objectives.
@@ V-G Lessons Learned
Additionally, our study delves into the dichotomy of individual versus collective efficacy, as reflected in ensemble learning’s portrayal of social individualism and collectivism. This also relates to the HRL’s directive versus autonomy debate, reflected in the balance of control and independence seen in society. Developing ML models that balance resilience and adaptability, while retaining specialization is challenging. The apparent synergies between ML and social dynamics beckon a future ripe with cross-disciplinary cooperation.
@@ VI-A1 Mutual Assistance
Figure 23: Illustration of Edge-edge Distributed Inference.
VI-A1a Edge-edge Distributed Inference
@@
  • Total Inference Latency Minimization: Edges accelerate inferences collaboratively, avoiding central delays. Just as society members possess varied abilities, edges too differ in resources. Addressing this involves task allocation and workload balancing. Works such as Wan et al. [163] and Mohammed et al. [164] delve into this, proposing DL-based strategies that maximize resource utilization in EI.
  • Total Cost Minimization: Beyond speed, holistic costs, such as energy and communication overheads, reflect sustainable social cooperation goals. The focus here is on achieving a balance and distributing tasks in a manner that minimizes collective costs. Innovations like those proposed by Fan et al. [165] have ventured into this area, optimizing edge collaborations for minimal costs.
@@
Edge-edge distributed inference, seen through the lens of mutual assistance, boasts several merits: 1) Privacy and Security. The communal edge approach safeguards sensitive data, echoing the social emphasis on individual privacy. 2) Scalability and Flexibility. Reflecting the adaptability of social cooperation, this method is inherently scalable and accommodative of diverse edges. 3) Low Latency. Emphasizing timely responses, the method caters to real-time applications, paralleling societies’ emphasis on immediacy in collaboration.
@@
Figure 24: Illustration of Device-device Collaborative DNN Inference.
@@
Device-device collaborative inference in EI highlights the combined efforts of devices, as illustrated in Fig. 24. Just as individuals in a society pool their strengths and resources to overcome challenges that are beyond their capability alone, devices, particularly those with limited computational capacities, can collaborate, exchange information, and process data jointly.
  • Total Inference Latency Minimization: This aspect involves the strategic sharing of computational tasks, akin to communal resource allocation, aimed at curtailing overall inference delays. Especially critical for UAVs and other latency-sensitive applications, this strategy involves optimizing task distribution across devices, even when faced with inherent resource limitations [166]. The objective remains to prevent resource overloading on a single device, ensuring harmonious operations [167].
  • Total Cost Minimization: The collaborative EI also extends to encompass cost-effectiveness, where devices, particularly those limited by battery life and storage, operate under a paradigm of energy efficiency. This approach necessitates a balance between energy expenditure for inference operations and communication, as highlighted in studies by Yang et al. [168] and Zeng et al. [169].
The goal is to optimally allocate tasks to devices to minimize both inference latency and energy consumption. The SL function Eq. (1) for device-device collaborative inference is $F([\text{devices}],[\text{mutual assistance}],\varnothing,\text{inference},\text{tasks})$. The optimization objective is:
@@
However, it also introduces challenges: 1) Privacy concerns: Increased inter-device communication potentially elevates the risk of data breaches, a dilemma paralleling privacy concerns in human societies. 2) Synchronization among heterogeneous devices: Variances in devices may complicate synchronization, an issue analogous to aligning individuals with diverse abilities in a society. 3) Task agility: The necessity for rapid adaptability in task management, considering variables like energy and computation, reflects the dynamism inherent in social collaborations.
Figure 25: Illustration of Edge-device Collaborative DNN Inference.
Communication Optimization in Edge-edge Distributed Inference: Edge-edge distributed inference heavily relies on efficient communication among edge devices to achieve optimal performance and minimize latency. One key challenge is to optimize the communication network topology and routing strategies to ensure fast and reliable information exchange. This can be formulated as a joint inference and communication optimization problem, where the objective is to minimize the total inference latency while satisfying the communication bandwidth and reliability constraints. Techniques such as network coding and cooperative communication can be applied to improve the communication efficiency and robustness in edge-edge distributed inference scenarios [170].
@@ VI-B1 Unidirectional Cooperation - Assistance
In EI during inference, certain computing layers provide insights or intermediary results to speed up others’ processes. The aided entity reaps the advantages while the helper plays a supporting role. Unlike training, where knowledge transfer is paramount, in EI’s inference phase, the focus shifts to utilizing specific layers or devices to minimize latency and expedite decisions, mirroring how society leverages resources for swift outcomes.
VI-B1a Edge-device Cooperative Inference
As depicted in Fig. 25, in edge-device cooperative inference, entities with greater resources offer support to those in need. Different computing layers collaborate to optimize DNN inference processes. The categorization of cooperative efforts can be understood through distinct performance optimization objectives:
  • Total Inference Latency Minimization: In EI, minimizing inference latency is analogous to the rapid social response to challenges. Methods like the DINA framework [164] segment DNNs for optimal task distribution, a strategy inspired by social cooperation and matching theory [171]. This emphasis on collaborative problem-solving for reduced response times is also championed in studies by Chen et al. [172] and Li et al. [173].
  • Inference Accuracy Maximization: Beyond speed, the quality of assistance is vital, reflected in EI’s pursuit of enhanced inference accuracy. Hanyao et al. [174] highlight this with a system that leverages cooperative inference to maintain accuracy under challenging conditions, mirroring the social drive for collective excellence. Similar themes are explored by Yun et al. [175] and Kong et al. [176], emphasizing how shared resources and knowledge lead to better outcomes.
  • Total Cost Minimization: Reflecting the social imperative for efficient resource use, EI strives for cost-effective operations. Fang et al. [177] demonstrate this through DNN filter pruning, optimizing resource utilization. This balance is also the focus of research by Du et al. [178] and Dai et al. [179], which mirror social efforts in optimal allocation.
@@
Edge-device collaborative DNN inference brings about: 1) Lower latency, emulating swift social assistance. 2) Lower cost, reminiscent of social efforts to optimize resource utilization. Despite the merits, as with social assistance, challenges persist. Real-time performance limitations mirror social constraints, emphasizing that while edge-device collaboration offers a fresh paradigm, its evolution is an extension, much like how social structures evolve over time.
Figure 26: Illustration of Cloud-device Collaborative DNN Inference.
Communication-Efficient Edge-Device Cooperative Inference: Edge-device cooperative inference involves frequent communication between edge servers and end devices, which can be a bottleneck in resource-constrained environments. To address this challenge, various communication-efficient strategies have been proposed, such as model compression, quantization, and selective transmission of intermediate features [180]. These techniques aim to reduce the amount of data exchanged between edge servers and devices while maintaining inference accuracy. Additionally, the design of communication-efficient protocols and scheduling algorithms is crucial for optimizing the overall performance of edge-device cooperative inference systems.
@@
Fig. 26 illustrates the collaborative DNN inference within EI, showcasing the dynamic between the cloud and the device that mirrors the social norm of unidirectional assistance. Here, the cloud, rich in resources, empowers devices to enhance their operational efficacy. This interplay is dissected further through the lens of performance optimization objectives:
  • Total Inference Latency Minimization: Vital for real-time applications such as autonomous driving and video surveillance, the balance between accuracy and latency is non-negotiable. Emmons et al. [181] highlight the merit of synergy between cloud and device. The introduction of complex DNN structures, such as GoogleNet [182] and ResNet [183], complicates the task of partitioning. This challenge has been addressed by Hu et al. [184] through the development of an adaptive scheme. Further research by Duan et al. [185] and Jeong et al. [186] emphasizes parallel DNN computation as a solution for latency.
@@
where $L=a\cdot L_{c}+(1-a)\cdot L_{d}$ is the total inference latency, $E=a\cdot E_{c}+(1-a)\cdot E_{d}$ is the total energy consumption, $a$ is the partition point of the DNN model, $L_{c}$ and $L_{d}$ are the inference latency on the cloud and device, respectively, $E_{c}$ and $E_{d}$ are the energy consumption on the cloud and device, respectively, and $\alpha$ and $\beta$ are the weights for latency and energy, respectively.
Figure 27: Illustration of Cloud-edge-device Collaborative DNN inference.
The parallels between assistance and cloud-device collaboration in inference are clear, offering: 1) Lowered inference latency, through strategic model partitioning and task allocation. 2) Reduced energy consumption, by optimizing communication and computation resources. Yet, challenges arise in: 1) Maintaining performance in unstable conditions, where rapid mobility or communication disruptions present hurdles. 2) Navigating the physical disconnect between cloud and device, limiting application scope. This comparison accentuates the essential synergy in achieving superior results, both in human social structures and in EI.
@@
Cloud-edge-device collaboration involves specific components (cloud, edge, or device) assisting others without expecting immediate reciprocal benefits. This collaborative model branches into three distinct categories, each focusing on enhancing various performance aspects, as shown in Fig. 27.
  • To address latency issues in collaborative DNN inference systems, Ren et al. [191] have formulated a distributed DNN computing partition. This setup coordinates computations across heterogeneous devices to boost inference efficiency. Further contributions by Xue et al. [192] focus on offloading and distributed DNN deployment, providing beneficial support to the whole system in terms of latency reduction.
  • Total Cost Minimization: Within this collaboration, cost reduction is integral, akin to a form of aid that elevates system functionality. Lin et al. [193] propose an adaptive optimization algorithm, reducing system offloading costs significantly. In a similar vein, strategies by Teerapittayanon et al. [194] and Liu and Wang et al. [195] develop approaches that strategically map parts of a DNN across the computing hierarchy, optimizing overall costs associated with communication and inference processes.
  • Failure-Resilient Distributed Model: In situations of component failures, the concept of assistance becomes paramount to ensure system robustness. The importance of assistance is portrayed by Yousefpour et al. [196], where skipping hyperconnections in distributed DNNs provide a safeguarding mechanism, ensuring continuity in the face of physical node failures.
@@
where $M$ is the entire DNN model, $M_{c}$, $M_{e}$, and $M_{d}$ are the partitions of the model executed on the cloud, edge, and device, respectively. $L(M_{x})$, $C(M_{x})$, and $F(M_{x})$ are the latency, cost, and failure resilience of executing partition $M_{x}$ on the corresponding platform, respectively. $\alpha,\beta,\gamma,\delta,\xi,\pi,\epsilon,\zeta,\eta$ are the weights for the respective terms.
Figure 28: Illustration of One-way Decision Correlation.
Compared to the prior two inference paradigms, cloud-edge-device collaborative inference serves as a supplement to both cloud-device and edge-device collaborative inferences, offering several advantages: 1) Optimized resource utilization, through detailed model segmentation that integrates the full potential of cloud centers, edges, and devices. 2) Enhanced fault recovery capabilities, by establishing a robust distributed system model that improves sensor fusion, system resilience, and data confidentiality.
@@
  • Total Inference Latency Minimization: One-way decision correlation streamlines processes by transferring decisions from one layer to another, reducing communication load. Benefiting from upper decisions, lower layers can enhance decision speed and quality, streamlining decision-making and reducing inference latency. Cao et al. [198] present a framework where the cloud ‘big model’ assists edge ‘small models’ by differentiating case complexity, and optimizing task distribution.
  • Energy Consumption Minimization: Similarly, in the energy consumption area, the lower layers align their operations efficiently by receiving and capitalizing on decisions from an upper layer, mitigating unnecessary computational expenses. The layer that benefits from this transfer of decisions saves energy, mirroring the dynamic where one party assists another by offering resources or expertise, leading to optimized outcomes. Kosuru et al. [199] showcase this principle with their methods, providing energy-efficient solutions in their respective contexts.
@@
$D_{i+1}=f_{i+1}(D_{i}),\ L_{\text{total}}=\sum_{i}L_{i}(D_{i}),\ E_{\text{total}}=\sum_{i}E_{i}(D_{i}),$ (26)
where $D_{i}$ is the decision made at layer $i$, $f_{i}$ is the decision-making function at layer $i$, $L_{i}(D_{i})$ and $E_{i}(D_{i})$ are the latency and energy consumption of making decision $D_{i}$ at layer $i$, respectively, and $L_{\text{total}}$ and $E_{\text{total}}$ are the total latency and energy consumption, respectively.
This non-reciprocal decision transfer aligns with altruism, streamlines processes, reduces inference latency, enhances real-time decision-making for time-sensitive tasks, and conserves energy by optimizing resource use and eliminating unnecessary computations.

### v2

@@ V-F1 Unidirectional Competition - Compulsory
Compulsory interaction is characterized by the dominance of one group and the subsequent compliance of another. Often, this stems from structural hierarchies, such as in workplaces where managers direct employees or in social constructs where certain groups have influence over others. In ML, this interaction can be observed in hierarchical structures where higher-level algorithms or protocols dictate the functioning or constraints of those at lower levels. By understanding this unidirectional competition, researchers can harness the efficiency and directionality it offers while being mindful of the inherent constraints and potential lack of flexibility it introduces.
@@
HRL mirrors unidirectional compulsory interactions where a dominant group or layer prescribes rules or guidelines for another group, mandating compliance. This structure is visually represented in Fig. 21. The essence of compulsory is exhibited as the superiors set overarching goals, while the subordinate layers are obligated to adhere. Dayan et al. [156] offer a quintessential example with a system partitioned into managers and sub-managers. Governed by principles like reward hiding and information hiding, this structure encapsulates unidirectional competition, with each layer, while autonomous to some extent, invariably remaining in contact with its superiors.
Figure 21: Illustration of hierarchical reinforcement learning in socialized training.
@@
where $H$ is the main function, $s$ is the state, $\pi^{G}$ and $\pi^{g}$ are the high-level and low-level policies, $\text{Transit}(\cdot)$ is the transition function, and $R$ is the reward function. Such hierarchical frameworks not only streamline learning but also afford the assimilation of prior knowledge. As in social structures, HRL embodies the directives of one group to the behaviour of another. The directives from higher layers, while potentially restrictive, shape the trajectory of the entire system, ensuring alignment with overarching objectives [157].
@@ V-F2 Bidirectional Competition - Struggle
The concept of struggle in social dynamics, when entities compete for shared resources, power, or influence, is reflected in adversarial learning techniques. In these techniques, two components challenge each other, and it’s through this continuous push and pull that the model evolves, yielding improved outcomes.
@@
V-F2a Adversarial Learning
In adversarial learning, particularly notable in GANs [158], there is a struggle between two components: the generator and the discriminator, as illustrated in Fig. 22. The generator crafts authentic samples to deceive the discriminator. Conversely, the discriminator endeavors to discern real data from the counterfeits created by the generator. However, challenges like instability in the training process and homogeneous sample generation persist. Original GANs often produced uncontrollable image variations from random noise.
[…]
$\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))],$
Solutions emerged in the form of conditional GANs [159], introducing conditional data and Deep Convolution GCNs (DCGANs) [160], swapping fully connected layers for convolutional ones. The quest for stability birthed the Wasserstein GAN [161] with its novel loss function. Drawing parallels to social struggle, adversarial learning represents inter-layer conflicts. This dynamic framework extends beyond image generation, spanning areas like RL [162] and natural language processing (NLP) [163].
Communication-aware Adversarial Learning: Adversarial learning techniques, such as GANs, have shown remarkable success in various applications. However, in a distributed setting, the communication between the generator and discriminator networks can be costly and may impact the overall performance. Communication-aware adversarial learning techniques have been proposed to address this issue by reducing communication overhead while maintaining the quality of the generated samples. These methods may involve strategies such as gradient compression, selective parameter sharing, and decentralized architectures for adversarial learning. The design of communication-efficient adversarial learning systems often requires considering the trade-off between communication cost and the stability and convergence of the training process [164].
@@
Decentralized Learning [89, 23, 90, 91, 92] Peer-to-peer collaboration among devices with limited resources. Efficient resource utilization, improved privacy, and robustness against single points of failure. Increased communication overhead, the potential for slower convergence.
@@
Meta Learning [135, 136] Applications requiring fast adaptation to new tasks or environments. Enables rapid learning and adaptation to new tasks, improves model generalization. Increased computational complexity, challenges in defining appropriate meta learning objectives.
@@ V-G Lessons Learned
Additionally, our study delves into the dichotomy of individual versus collective efficiency, as reflected in ensemble learning’s portrayal of social individualism and collectivism. This also relates to the HRL’s directive versus autonomy debate, reflected in the balance of control and independence seen in society. Developing ML models that balance resilience and adaptability while retaining specialization is challenging. The apparent synergy between ML and social dynamics signals a future maturity for cross-disciplinary cooperation.
Figure 23: Illustration of edge-edge collaborative inference.
@@ VI-A1 Mutual Assistance
VI-A1a Edge-edge Collaborative Inference
@@
  • Total Inference Latency Minimization: Edges accelerate inferences collaboratively, avoiding central delays. Just as members of society possess varied abilities, edge resources also are different. Addressing this involves task allocation and workload balancing. Studies such as Wan et al. [166] and Mohammed et al. [167] delve into this, proposing DL-based strategies that maximize resource utilization in EI.
  • Total Cost Minimization: Beyond speed, holistic costs, such as energy and communication overheads, reflect sustainable social cooperation goals. The focus here is on achieving a balance and distributing tasks in a manner that minimizes collective costs. Studies like those proposed by Fan et al. [168] have ventured into this area, optimizing edge collaborations for minimum costs.
@@
Figure 24: Illustration of device-device collaborative inference.
Edge-edge collaborative inference, seen through the lens of mutual assistance, boasts several merits: 1) Privacy and Security. The communal edge approach safeguards sensitive data, echoing the social emphasis on individual privacy. 2) Scalability and Flexibility. Reflecting the adaptability of social cooperation, this method is inherently scalable and accommodative of diverse edges. 3) Low Latency. Emphasizing timely responses, the method caters to real-time applications, paralleling societies’ emphasis on immediacy in collaboration.
@@
Communication Optimization in Edge-edge Collaborative Inference: Edge-edge collaborative inference heavily relies on efficient communication among edges to achieve optimal performance and minimize latency. One key challenge is to optimize the communication network topology and routing strategies to ensure fast and reliable information exchange. This can be formulated as a joint inference and communication optimization problem, aiming to minimize total inference latency while satisfying communication bandwidth and reliability constraints. Techniques such as network encoding and cooperative communication can be applied to improve the communication efficiency and robustness in edge-edge collaborative inference scenarios [169].
@@
Device-device collaborative inference in EI highlights the combined efforts of devices, as illustrated in Fig. 24. Just as individuals in a society pool their strengths and resources to overcome challenges that are beyond their capability alone, devices, particularly those with limited computational capacities. They can collaborate, exchange information, and process data jointly.
  • Total Inference Latency Minimization: This aspect involves the strategic sharing of computational tasks, akin to public resource allocation, aimed at reducing overall inference delays. Especially critical for UAVs and other latency-sensitive applications, this strategy involves optimizing task allocation across devices, even when faced with inherent resource limitations [170]. The objective remains to prevent resource overloading on a single device, ensuring harmonious operations [171].
  • Total Cost Minimization: The collaborative EI also extends to encompass cost-effectiveness, where devices, particularly those limited by battery life and storage, operate under a paradigm of energy efficiency. This approach necessitates a balance between energy consumption for inference operations and communication, as highlighted in studies by Yang et al. [172] and Zeng et al. [173].
The objective is to optimally allocate tasks to devices to minimize both inference latency and energy consumption. The SL function Eq. (1) for device-device collaborative inference is $F([\text{devices}],[\text{mutual assistance}],\varnothing,\text{inference},\text{tasks})$. The optimization objective is:
@@
Figure 25: Illustration of edge-device collaborative inference.
@@
However, it also introduces challenges: 1) Privacy concerns: Increased inter-device communication potentially elevates the risk of data leakage, a dilemma paralleling privacy concerns in human societies. 2) Synchronization among heterogeneous devices: Variances in devices may complicate synchronization, an issue analogous to aligning individuals with diverse abilities in a society. 3) Task adaption: The necessity for rapid adaptability in task management, considering variables like energy and computation, reflects the dynamism inherent in social collaborations.
@@ VI-B1 Unidirectional Cooperation - Assistance
In EI during inference, certain computing layers provide insights or intermediary results to speed up others’ processes. The aided entities gain the advantages while the helpers play a supporting role. Unlike training, where knowledge transfer is paramount, in EI’s inference phase, the focus shifts to utilizing specific layers or devices to minimize latency and expedite speed up decision-making, mirroring how society leverages resources for swift outcomes.
VI-B1a Edge-device Collaborative Inference
As depicted in Fig. 25, in edge-device collaborative inference, entities with greater resources offer support to those in need. Different computing layers collaborate to optimize DNN inference processes. The categorization of assistant efforts can be understood through distinct performance optimization objectives:
  • Total Inference Latency Minimization: In EI, minimizing inference latency is analogous to the rapid social response to challenges. Methods like the DINA framework [167] segment DNNs for optimal task allocation, a strategy inspired by social cooperation and matching theory [174]. This emphasis on collaborative problem-solving for reduced response times is also proposed in studies by Chen et al. [175] and Li et al. [176].
  • Inference Accuracy Maximization: Beyond speed, the quality of assistance is vital, reflected in EI’s pursuit of enhanced inference accuracy. Hanyao et al. [177] highlight this with a system that leverages collaborative inference to maintain accuracy under challenging conditions. Similar studies are explored by Yun et al. [178] and Kong et al. [179], emphasizing how shared resources and knowledge can lead to better performance.
  • Total Cost Minimization: Reflecting the social requirement for efficient resource use, EI strives for cost-effective operations. Fang et al. [180] demonstrate this through DNN filter pruning, optimizing resource utilization. This balance is also the focus of research by Du et al. [181] and Dai et al. [182], which mirror social efforts in optimal allocation.
@@
Edge-device collaborative DNN inference brings about: 1) Lower latency, akin to rapid social assistance. 2) Lower cost, analogous to social efforts aimed at optimizing resource utilization. Despite these advantages, challenges analogous to those in social assistance persist. Limitations in real-time performance reflect social constraints, underscoring that while it introduces a novel paradigm, its evolution is progressive, similar to the evolution of social structures over time.
Figure 26: Illustration of cloud-device collaborative inference.
Communication-efficient Edge-device Collaborative Inference: Edge-device collaborative inference involves frequent communication between ESs and devices, which can be a bottleneck in resource-constrained environments. To address this challenge, various communication-efficient strategies have been proposed, such as model compression, quantization, and selective transmission of intermediate features [183]. These techniques aim to reduce the amount of data exchanged between ESs and devices while maintaining inference accuracy. Additionally, the design of communication-efficient protocols and scheduling algorithms is crucial for optimizing the overall performance of edge-device collaborative inference systems.
@@
Fig. 26 illustrates the collaborative DNN inference within EI, showcasing the dynamic between the cloud and the device that mirrors the social norm of unidirectional assistance. Here, the cloud, rich in resources, empowers devices to enhance their operational efficiency. This interplay is discussed further through the lens of performance optimization objectives:
  • Total Inference Latency Minimization: Vital for real-time applications such as autonomous driving and video surveillance, the balance between accuracy and latency is non-negotiable. Emmons et al. [184] highlight the merit of synergy between cloud and devices. The introduction of complex DNN structures, such as GoogleNet [185] and ResNet [186], complicates the model partitioning. This challenge has been addressed by Hu et al. [187] through the development of an adaptive scheme. Further research by Duan et al. [188] and Jeong et al. [189] emphasizes parallel DNN computation as a solution for latency.
@@
where $L=aL_{c}+(1-a)L_{d}$ is the total latency, $E=aE_{c}+(1-a)E_{d}$ is the total energy consumption, $a$ is the partition point of DNN, $L_{c}$ and $L_{d}$ are the latency on the cloud and device, respectively, $E_{c}$ and $E_{d}$ are energy consumption on the cloud and device, respectively, and $\alpha$ and $\beta$ are weights for latency and energy consumption, respectively.
Figure 27: Illustration of cloud-edge-device collaborative inference.
The comparison between assistance and cloud-device collaboration in inference clearly demonstrates similar advantages, such as: 1) Lowered inference latency, achieved through strategic model partitioning and task allocation. 2) Reduced energy consumption, by optimizing the use of communication and computation resources. However, this collaboration also faces distinct challenges: 1) Maintaining performance in unstable conditions, where factors like rapid mobility or communication disruptions can present significant hurdles. 2) Navigating the physical disconnect between cloud and devices, which can limit the scope of applications. This comparison underscores the essential synergy required to achieve superior results in both human society and EI.
@@
Cloud-edge-device collaboration involves specific components (cloud, edge, or device) assisting others without expecting immediate reciprocal benefits. This collaboration branches into three distinct categories, each focusing on enhancing various performance aspects, as shown in Fig. 27.
  • Total Inference Latency Minimization: To address latency issues in collaborative DNN inference systems, Ren et al. [194] formulate a distributed DNN partition. This setup coordinates computations across heterogeneous entities to boost inference efficiency. Further contributions by Xue et al. [195] focus on offloading and distributed DNN deployment, providing beneficial support to the whole system in terms of latency reduction.
  • Total Cost Minimization: Within this collaboration, cost reduction is integral, akin to a form of aid that elevates system functionality. Lin et al. [196] propose an adaptive optimization algorithm, reducing system offloading costs significantly. Similarly, strategies by Teerapittayanon et al. [197] and Liu and Wang et al. [198] develop approaches that strategically map parts of a DNN across the computing hierarchy, optimizing overall costs associated with communication and inference processes.
  • Failure-Resilient Distributed Model: In situations of component failures, the concept of assistance becomes paramount to ensure system robustness. The importance of assistance is portrayed by Yousefpour et al. [199], where skipping hyperconnections in distributed DNNs provides a safeguarding mechanism, ensuring continuity in the face of physical node failures.
@@
where $M$ is the entire DNN, $M_{c}$, $M_{e}$, and $M_{d}$ are the partitions of the model executed on the cloud, edge, and device, respectively. $L(M_{x})$, $C(M_{x})$, and $F(M_{x})$ are the latency, cost, and failure resilience of executing partition $M_{x}$ on the corresponding platform, respectively. $\alpha,\beta,\gamma,\delta,\xi,\pi,\epsilon,\zeta,\eta$ are the weights for the respective terms.
Figure 28: Illustration of one-way decision correlation in socialized inference.
Compared to the prior two inference paradigms, cloud-edge-device collaborative inference serves as a supplement to both cloud-device and edge-device collaborative inferences, offering several advantages: 1) Optimized resource utilization, through detailed model segmentation that integrates the full potential of clouds, edges, and devices. 2) Enhanced fault recovery capabilities, by establishing a robust distributed system model that improves sensor fusion, system resilience, and data confidentiality.
@@
  • Total Inference Latency Minimization: One-way decision correlation streamlines processes by transferring decisions from one layer to another, reducing communication load. Benefiting from upper decisions, lower layers can enhance decision speed and quality, streamlining decision-making and reducing inference latency. Cao et al. [201] present a framework where the cloud ‘big model’ assists edge ‘small models’ by differentiating case complexity, and optimizing task allocation.
  • Energy Consumption Minimization: Similarly, in the energy consumption area, the lower layers align their operations efficiently by receiving and capitalizing on decisions from an upper layer, mitigating unnecessary computational costs. The layer that benefits from this transfer of decisions saves energy, mirroring the dynamic where one party assists another by offering resources or expertise, leading to optimized outcomes. Kosuru et al. [202] showcase this principle within their methods, providing energy-efficient solutions in their respective contexts.
@@
min $\quad\sum_{i}L_{i}(D_{i})+\sum_{i}E_{i}(D_{i}),\ D_{i+1}=f_{i+1}(D_{i}),$ (26)
where $D_{i}$ is the decision made at layer $i$, $f_{i}$ is the decision-making function at layer $i$, $L_{i}(D_{i})$ and $E_{i}(D_{i})$ are the latency and energy consumption of making decision $D_{i}$ at layer $i$, respectively. This non-reciprocal decision transfer aligns with altruism, streamlines processes, reduces inference latency, enhances real-time decision-making for time-sensitive tasks, and conserves energy by optimizing resource use and eliminating unnecessary computations.